write a equation of a given line that passes parrella through the given point? x+y=3 (5,4)
Answer by stanbon(75887)

If x+y=3 then y=-x+3.
Therefore the given slope is m=-1
The point gives yoy x=5,y=4
The equation of the line is y=mx+b
Substitute the values to get:
4=(-1)(5)+b
So, b= 9
Therefore the equation you want is y = -x+9
cheers,
Stan H.